# Identifying Information from Observations with Uncertainty and Novelty

### v1

@@ 4.2 Identifying Information from Direct Observations
We now prove properties of identification and sample complexity when the set of known parameters $\mathbf{\Theta}$ is also the sample space of the observed string $\theta$. In this case, there is no hidden or latent state space as the parameter space is the observed sample space. Note that $\mathbf{\Theta}$ is a mathematical set where all the strings are unique, however in order to compute set membership, the comparison of the index-aligned symbols in the strings requires there to be a chosen order of comparison. Every mathematical set that is computable is fully identifiable when the strings in the set are what is observed, as we prove here and later, but they may not be fully identifiable when what is observed is the output of a computable function of the set’s strings. Given computation is our concern, these sets will often be manipulated as sequences.
@@
3:     $j\leftarrow 1$triangleright The index or number of elements checked to identify
@@
$\exists(\vec{\psi\mkern-3.0mu}\mkern 3.0mu\in\mathbf{\Theta}):\big{(}\forall(j\in\mathbb{Z}^{+}):\vec{\psi\mkern-3.0mu}\mkern 3.0mu_{j}=\vec{\theta\mkern-3.0mu}\mkern 3.0mu_{j}\big{)}$
@@
$\forall(\vec{\psi\mkern-3.0mu}\mkern 3.0mu\in\mathbf{\Theta}):\big{(}\exists(j\in\mathbb{Z}^{+}):\vec{\psi\mkern-3.0mu}\mkern 3.0mu_{j}\neq\vec{\theta\mkern-3.0mu}\mkern 3.0mu_{j}\big{)}$
@@
  • there is an accepted encoding of the parameters and the observations to a shared string representation space
@@ 4.3.1 Finite Information Spread Over Infinite Observations
($\implies$) Given the observer knows everything but the symbols at each index and symbols’ corresponding Bernoulli distributions, this may be treated as learning which component distribution is almost surely used at each index of $\theta$. Treating this as $\ell(\theta)$ different empirical processes updated as one cycles through the indices of $\theta$, we know by the Glivenko-Cantelli theorem that the empirical processes will asymptotically converge almost surely to their true distributions used at each index as the observations goes to infinite. The different distributions for the symbols are almost surely identifiable from one another as $\lim_{t\rightarrow\infty}\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1}$ and thus the original bits of $\theta$ are then identified.
@@ 4.4 Partial Identification and the Distribution of the Sample Complexity
(Partial Identification of an I.I.D. Process with Probability $p$) Let the computable hypothesis set be $\mathbf{\Theta}:|\mathbf{\Theta}|<\infty,\forall(\theta\in\mathbf{\Theta}):(\ell(\theta)<\infty)$, the fixed identifiable parameter $\phi\in\mathbf{\Theta}$, $\lambda_{S}(\phi,t,Z)=\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1}$ be the observations from an i.i.d. process $\vec{x\mkern-3.0mu}\mkern 3.0mu\sim X$, and the prior distribution be such that $\forall(\theta\in\mathbf{\Theta}):P(\vec{\Theta\mkern-3.0mu}\mkern 3.0mu_{0}=\theta)<p$, where $10<p<$ 1 is the desired minimum probability that a model $\theta$ generated the observations. Then, the sample complexity $\mathbf{i}(\vec{\mathbf{X}\mkern-3.0mu}\mkern 3.0mu^{t}_{1},\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1},p,0,0,0)$ forms a random variable over $\mathbb{Z}^{+}$ that is determined by its moments, which are surely finite. The expected sample complexity is
@@
So far, we have relied upon $\lambda_{S}(\theta,t,Z)=\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1}$ to compute $t$ samples from an arbitrary i.i.d. random variable. To extend our scope to a stationary ergodic process with finite $L$ memory we introduce the computation of its transition function $T$ by a Turing Machine with the description of $T$ denoted as $\lambda_{T}(\delta,\vec{x\mkern-3.0mu}\mkern 3.0mu^{-L}_{t})=\theta_{t}$, where $\delta$ is a finite description of a mapping of at most $2^{L}$ length strings $\vec{x\mkern-3.0mu}\mkern 3.0mu^{-L}_{t}$ to the unique descriptions of i.i.d. variables. $\lambda_{S}(\theta_{t},1,Z)=x_{t}$ then computes the sample at each time-step. To initiate computation of the process and to account for the $L$ prior state dependency, either $x^{-L}_{1}$ is provided along with $\delta$, or $\delta$ includes a means of determining $x^{-L}_{1}$. In the case of a randomized initial state, $\delta$ may contain another distribution for sampling which of the $2^{L}$ possible random variables are sampled from for the initial $L$ symbols . If $\delta$ describes the joint probability distribution of $P(\vec{X\mkern-3.0mu}\mkern 3.0mu^{L+1}=\vec{x\mkern-3.0mu}\mkern 3.0mu^{L+1})$, then given any $L+1$ length string with missing symbols, the probable distributions for the missing samples can then be informed by $\delta$ and even sampled to fill in the missing samples, although we primarily consider the conditional $P(\vec{X\mkern-3.0mu}\mkern 3.0mu_{t}=\vec{x\mkern-3.0mu}\mkern 3.0mu_{t}|\vec{X\mkern-3.0mu}\mkern 3.0mu^{-L}_{t}=\vec{x\mkern-3.0mu}\mkern 3.0mu^{-L}_{t})$ for computing the next output. After the output is populated with $L$ symbols, the process simply reads the $L$ prior output to determine the next $\theta$ to sample from, and repeats. This computation allows for any dependency on the prior $L$ symbols and also includes any deterministic process with $L$ memory. With this computation in mind, we now show how such processes can be identified by predictors that can learn processes with less than or equal to $L$ memory.
@@ 4.4.2 Probable Exhaustive Falsification: Probable Novelty
A misspecified hypothesis set is when $\phi\notin\mathbf{\Theta}$ (Ghosal and Vaart, 2017, Ch. 8.5) and is known as “agnostic learning” in statistical learning (Shalev-Shwartz and Ben-David, 2014, Def. 3.3). The minimum error is only zero if $\phi\in\mathbf{\Theta}$, which is referred to as “realizable learning” in statistical learning (Shalev-Shwartz and Ben-David, 2014, Def. 2.1). The concept of a misspecified hypothesis set raises the question of “How can one detect that the evidence is more likely generated from an unknown model rather than any of the known models?” This is not a new question. This is the same question asked by “goodness of fit” tests of a single known model given the evidence, rather than comparing two or more models against each another as per Bayes factors, the Bayes information criterion, odds ratios, or maximum likelihood. The above theorems already consider the case of comparison of multiple known models. In the end, this problem is relatively more underspecified than the comparison of known models because we only know our considered models, and so can only observe how probable the observations are given those models. Following the principle of maximum likelihood (Stigler, 1997, 2007), one can subjectively decide how likely such an observed sequence need be given the known model, or the posterior mixture of known models. This subjective decision is the same as determining the desired certainty in a confidence or credible interval, and in our case will determine the typical sets for the sample complexity in Definition 4.1.7.
@@
(Sample Complexity of Probable Exhaustive Falsification) Let the computable hypothesis set be $\mathbf{\Theta}:|\mathbf{\Theta}|<\infty,\forall(\theta\in\mathbf{\Theta}):(\ell(\theta)<\infty)$, $\lambda_{S}(\mathbf{\Theta},Z)=\vec{\mathbf{X}\mkern-3.0mu}\mkern 3.0mu^{*}_{1}$ be a set of stationary ergodic processes, the fixed finite process description $\phi\notin\mathbf{\Theta}$, and $\lambda_{S}(\phi,Z)=\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1}$ be the observations of a stationary ergodic process. If $q>0$, then $\mathbf{i}(\vec{\mathbf{X}\mkern-3.0mu}\mkern 3.0mu^{*}_{1},\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1},1,q,0,0)$ forms a distribution over $Z^{+}$ that is determined by its moments, which are surely finite.
@@
The typical set’s complement defines the atypical set, which is the union of a pair of disjoint sets, i.e., the atypically improbable set and atypically probable set of observations for the given model. To determine observations to be atypically improbable is to indicate that on average over the possible samples there exists a better model of those observations for which they are more probable. This informs to change the average direction of probable and improbable events. For atypically probable, there is a better model where the observations that are probable for this model are more probable and, similarly by conservation of probability, those that are improbable are more improbable. This indicates on average that the probability distribution has the correct ordering of the probability of events, but needs to increase the contrast between the improbable and probable events. When considering multiple known models’ atypical sets and weighting them by a probability distribution, this informs how to update relative weighting and recovers the Bayesian estimation approach. When all known models indicate an atypical set, they serve as points in the probability space to inform the direction of the unknown ideal model relative to them, which is more informative than only knowing that the ideal is probably none of the known models.
@@ 5 Conclusion
In this work, we unified the identification problem across various fields, including theory of computation, asymptotic statistics, and Bayesian probably approximately correct learning from statistical learning theory. We covered different cases of computing the indicator function within finite observations, from deterministic scenarios to ergodic stationary stochastic processes. We demonstrated the properties of the sample complexity in these scenarios and proved that in the PAC-Bayes case, the sample complexity is determined by its moments, all of which are finite. This means that it is asymptotically bounded by an exponential decay, and may be computed to any desired accuracy based on the chosen prior. In each of these cases, we addressed what novelty is, and showed that it is identified by (probable) exhaustive falsification of the known models given the observations. From this work, we have more complete answers to the questions raised in the abstract. Information is determined by the number of distinct states and their relationships that can be expressed in a language, such as a programming language or a formal language. The information provided by the observations is relative to the observation space and the set of hypotheses considered, whose superset is determined by the chosen formal language, programming language, and resource limitations, such as memory space limitations. The hypothesis set and its representation within your language determine how many observations remain to verify that an element in the set corresponds to the observations or exhaustively falsify the hypothesis set, which is to identify novelty. Although our formalization of identifying information with probability and error for a fixed finite hypothesis set covers the identification of novelty, this does not address the learning of a new hypothesis from observations, as in the case of a finitely growing or changing hypothesis set. We have established the necessary foundation for such a work to be built upon, which will be explored in the future.

### v2

@@ 4.2 Identifying Information from Direct Observations
We now prove properties of identification and sample complexity when the set of known hypotheses $\mathbf{\Theta}$ is also the sample space of the observed string $\theta$. In this case, there is no hidden or latent state space as the hypothesis space is the observed sample space. Note that $\mathbf{\Theta}$ is a mathematical set where all the strings are unique, however in order to compute set membership, the comparison of the index-aligned symbols in the strings requires there to be a chosen order of comparison. Every mathematical set that is computable is fully identifiable when the strings in the set are what is observed, as we prove here and later, but they may not be fully identifiable when what is observed is the output of a computable function of the set’s strings. Given computation is our concern, these sets will often be manipulated as sequences.
@@
3:     $j\leftarrow 1$triangleright The index or number of elements checked
@@
$\exists(\vec{\psi\mkern-3.0mu}\mkern 3.0mu\in\mathbf{\Theta}):\big{(}\forall(j\in\{1,2,...,\ell(\vec{\psi\mkern-3.0mu}\mkern 3.0mu)\}):\vec{\psi\mkern-3.0mu}\mkern 3.0mu_{j}=\vec{\theta\mkern-3.0mu}\mkern 3.0mu_{j}\big{)}$
@@
$\forall(\vec{\psi\mkern-3.0mu}\mkern 3.0mu\in\mathbf{\Theta}):\big{(}\exists(j\in\{1,2,...,\ell(\vec{\psi\mkern-3.0mu}\mkern 3.0mu)\}):\vec{\psi\mkern-3.0mu}\mkern 3.0mu_{j}\neq\vec{\theta\mkern-3.0mu}\mkern 3.0mu_{j}\big{)}$
@@
  • there is an accepted encoding of the hypotheses’ descriptions and the observations to a shared string representation space
@@ 4.3.1 Finite Information Spread Over Infinite Observations
($\implies$) Given the observer knows everything but the symbols at each index and symbols’ corresponding Bernoulli distributions, this may be treated as learning which component distribution is almost surely used at each index of $\theta$. Treating this as $\ell(\theta)$ different empirical processes updated as one cycles through the indices of $\theta$, we know by the Glivenko-Cantelli theorem that the empirical processes will asymptotically converge almost surely to their true distributions used at each index as the observations goes to infinite. The different distributions for the symbols are almost surely identifiable from one another as $\lim_{t\rightarrow\infty}\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1}$ and thus the original bits of $\theta$ are also almost surely identified.
@@ 4.4 Partial Identification and the Distribution of the Sample Complexity
[…]
@@
So far, we have relied upon $\lambda_{S}(\theta,t,\vec{z\mkern-3.0mu}\mkern 3.0mu)=\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1}$ to compute $t$ samples from an arbitrary i.i.d. random variable. To extend our scope to a stationary ergodic process with finite $L$ memory we introduce the computation of its transition function $T$ by a Turing Machine with the description of $T$ denoted as $\lambda_{T}(\delta,\vec{x\mkern-3.0mu}\mkern 3.0mu^{-L}_{t})=\theta_{t}$, where $\delta:\mathrsfso{X}^{L}\mapsto\mathbf{\Theta}$ is finite mapping of at most $2^{L}$ length strings to descriptions of i.i.d. variables. $\lambda_{S}(\theta_{t},1,\vec{z})=x_{t}$ then computes the sample at each time-step. We use $\lambda_{TS}(\delta,\vec{x\mkern-3.0mu}\mkern 3.0mu^{-L}_{t},i,\vec{z})=\vec{x\mkern-3.0mu}\mkern 3.0mu^{i}_{t}$ to denote the chained alternating computations of $\lambda_{T}$ and $\lambda_{S}$ to compute an $i$ length sampling from a process whose description $\theta_{t}$ changes over the sampling conditioned on the moving window of $L$ prior samples.
To initiate computation of the process and to account for the $L$ prior state dependency, either $x^{-L}_{1}$ is provided along with $\delta$, or $\delta$ includes a means of determining $x^{-L}_{1}$. In the case of a randomized initial state, $\delta$ may contain another distribution for sampling which of the $2^{L}$ possible random variables are sampled from for the initial $L$ symbols . If $\delta$ describes the joint probability distribution of $P(\vec{X\mkern-3.0mu}\mkern 3.0mu^{L+1}=\vec{x\mkern-3.0mu}\mkern 3.0mu^{L+1})$, then given any $L+1$ length string with missing symbols, the probable distributions for the missing samples can then be informed by $\delta$ and even sampled to fill in the missing samples, although we primarily consider the conditional $P(\vec{X\mkern-3.0mu}\mkern 3.0mu_{t}=\vec{x\mkern-3.0mu}\mkern 3.0mu_{t}|\vec{X\mkern-3.0mu}\mkern 3.0mu^{-L}_{t}=\vec{x\mkern-3.0mu}\mkern 3.0mu^{-L}_{t})$ for computing the next output. After the output is populated with $L$ symbols, the process simply reads the $L$ prior output to determine the next $\theta$ to sample from, and repeats. This computation allows for any dependency on the prior $L$ symbols and also includes any deterministic process with $L$ memory. With this computation in mind, we now show how such processes can be identified by predictors that can learn processes with less than or equal to $L$ memory.
@@ 4.4.2 Probable Exhaustive Falsification: Probable Novelty
A misspecified hypothesis set is when $\phi\notin\mathbf{\Theta}$ (Ghosal and Vaart, 2017, Ch. 8.5) and is known in statistical learning as “agnostic learning” where the minimum error is greater than zero (Shalev-Shwartz and Ben-David, 2014, Def. 3.3). The minimum error is only zero if $\phi\in\mathbf{\Theta}$, which is referred to as “realizable learning” in statistical learning (Shalev-Shwartz and Ben-David, 2014, Def. 2.1). The concept of a misspecified hypothesis set raises the question of “How can one detect that the evidence is more likely generated from an unknown model rather than any of the known models?” This is not a new question. This is the same question asked by “goodness of fit” tests of a single known model given the evidence, rather than comparing two or more models against each another as per Bayes factors, the Bayes information criterion, odds ratios, or maximum likelihood. The above theorems already consider the case of comparison of multiple known models. In the end, this problem is relatively more underspecified than the comparison of known models because we only know our considered models, and so can only observe how probable the observations are given those models. Following the principle of maximum likelihood (Stigler, 1997, 2007), one can subjectively decide how likely such an observed sequence need be given the known model, or the posterior mixture of known models. This subjective decision is the same as determining the desired certainty in a confidence or credible interval, and in our case will determine the typical sets for the sample complexity in Definition 4.1.7.
@@
(Sample Complexity of Probable Exhaustive Falsification) Let the computable set of hypothesized latent states be $\mathbf{\Theta}:|\mathbf{\Theta}|<\infty,\forall(\theta\in\mathbf{\Theta}):(\ell(\theta)<\infty)$, $\mathbf{\Delta}\subseteq\mathrsfso{X}^{L}\times\mathbf{\Theta}:\{\delta\in\Delta:\mathrsfso{X}^{L}\mapsto\mathbf{\Theta}\}$, $\vec{z\mkern-3.0mu}\mkern 3.0mu\sim Z$ be the results of flipping a fair coin, $\lambda_{TS}(\mathbf{\Delta},\mathrsfso{X}^{L},*,Z)=\vec{\mathbf{X}\mkern-3.0mu}\mkern 3.0mu^{*}_{1}$ be a set of stationary ergodic processes, the fixed finite process description $\phi\notin\mathbf{\Delta}$, and $\lambda_{TS}(\phi,\vec{x\mkern-3.0mu}\mkern 3.0mu^{-L}_{1},t,\vec{z\mkern-3.0mu}\mkern 3.0mu)=\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1}$ be the observations of a stationary ergodic process. If $q>0$, then $\mathbf{i}(\vec{\mathbf{X}\mkern-3.0mu}\mkern 3.0mu^{*}_{1},\vec{x\mkern-3.0mu}\mkern 3.0mu^{t}_{1},1,q,0,0)$ forms a distribution over $\mathbb{Z}^{+}$ that is determined by its moments, which are surely finite.
@@
The typical set’s complement defines the atypical set, which is the union of a pair of disjoint sets, i.e., the atypically improbable set and atypically probable set of observations for the given model. The earlier Figure 3 visualizes an example’s typical and atypical sets over the hypothesis’ probability measure of the observations and shows how we allow for an undetermined set between the typical and atypical set of observations when $p>q$. To determine observations to be atypically improbable is to indicate that on average over the possible samples there exists a better model of those observations for which they are more probable. This informs to change the average direction of probable and improbable events. For atypically probable, there is a better model where the observations that are probable for this model are more probable and, similarly by conservation of probability, those that are improbable are more improbable. This indicates on average that the probability distribution has the correct ordering of the probability of events, but needs to increase the contrast between the improbable and probable events. When considering multiple known models’ atypical sets and weighting them by a probability distribution, this informs how to update relative weighting and recovers the Bayesian estimation approach. When all known models indicate an atypical set, they serve as points in the probability space to inform the direction of the unknown ideal model relative to them, which is more informative than only knowing that the ideal is probably none of the known models.
@@
Given the empirical process will not be able to measure within our accepted probability at least until it has observed $\vec{H\mkern-3.0mu}\mkern 3.0mu(\vec{X\mkern-3.0mu}\mkern 3.0mu)-q$ bits, there is a “warm up” period necessary for more reliable evidence that the hypotheses are probably exhaustively falsified. Figure 3(a) depicts this where the empirical process will estimate probability one for the single observation, which would put it well within one of the atypical sets for rejecting the hypothesis. Figure 3(a) is to visualize the sets clearly and demonstrate this important point that it does not make sense to compare the hypothesis to an undersampled representation as simpler hypotheses will favor smaller observation sequences due to fewer degrees of freedom required to be explained by evidence, i.e., the principle of parsiomony, popularly known as Occam’s razor (Charlesworth, 1955).
@@ 5 Conclusion
In this work, we explored the computation of hypothesis identification. Identifying novel information, as well as verifying hypotheses or their parts, whether in an absolute binary fashion or in degrees, enables learning the structure of the underlying data-generating process. Fundamentally, identification is required for a machine to learn. Identification also informs a predictor when its hypothesis set is misspecified, and thus in need of an update. We unified the identification problem across various fields, including the theory of computation, asymptotic statistics, and Bayesian probably approximately correct learning from statistical learning theory. We covered different cases of computing the indicator function within finite observations, from deterministic scenarios to ergodic stationary stochastic processes. We demonstrated the properties of the sample complexity in these scenarios and proved that in the PAC-Bayes case, the sample complexity is determined by its moments, all of which are finite. This means that it is asymptotically bounded by an exponential decay, and may be computed to any desired accuracy based on the chosen prior. In each of these cases, we addressed what novelty is and showed that it is identified by (probable) exhaustive falsification of the known models given the observations.
From this work, we have more complete answers to the questions raised in the abstract. Information is determined by the number of distinct states and their relationships that can be expressed in a language, such as a programming language or a formal language. The information provided by the observations is relative to the observation space and the set of hypotheses considered, whose superset is determined by the chosen formal language, programming language, and resource limitations, such as memory space limitations. The hypothesis set and its representation within your language determine how many observations remain to verify that an element in the set corresponds to the observations or exhaustively falsify the hypothesis set, which is to identify novelty. Although our formalization of identifying information with probability and error for a fixed finite hypothesis set covers the identification of novelty, this does not address the learning of a new hypothesis from observations, as in the case of a finitely growing or changing hypothesis set. We have established the necessary foundation for such a work to be built upon, which will be explored in the future.